OpenStudy (anonymous):

The table below represents a linear function f(x) and the equation represents a function g(x) x f(x) -1 -11 0 -1 1 9 g(x) g(x)=5x+1 Part A: Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x). Part B: Which function has a greater y-intercept? Justify your answer.

OpenStudy (anonymous):

I could answer this, first though, you're doing this on FLVS right?

OpenStudy (radar):

The x function is linear (problem says so) from the table develop the equation for the line. You are given three points, use any 2 and do it.

OpenStudy (anonymous):

Yes @LearningIsAwesome

OpenStudy (anonymous):

Copying and pasting questions from your school is against the Terms of Use you agreed to when signing up. Please close the question.

OpenStudy (anonymous):

okay first of all @LearningIsAwesome I just wanted help so that I can understand it better second almost half of the questions on this are from flvs so yeah

OpenStudy (anonymous):

That's okay, just don't copy and paste questions that's all :) bye

OpenStudy (anonymous):

so any way @radar do I just pick 2 from the graph and what exactly do I do with them

OpenStudy (anonymous):

@radar

OpenStudy (radar):

determine slope. This can be done by the formula slope m - change in y (delta y) divided by change in x (delta x)\[\Delta y \over \Delta x\] that is the formula for slope. Now in that table f(x) would be the y value, and of course x is the x value. Use the values provided by the problem.

OpenStudy (radar):

The change can be obtained by subtraction. Taking the difference.

OpenStudy (radar):

Yes, just pick two points and be consistent with your subtraction, other words you are subtracting one point from the other (P1 - P2 where the x =x, and y = f(x).

OpenStudy (radar):

Excuse me if I don't respond quickly, I am having a pizza for lunch.

OpenStudy (anonymous):

oh so I choose two points so say I choose -11 - (-1) and 1-0 and then divide them?

OpenStudy (radar):

Good, you have got the right idea, I will use the same points and we can compare our answers. -11 -(-1) = -10 and -1 - 0 =-1 So -10/-1 = 10 a positive 10. Did you get the same?

OpenStudy (anonymous):

yes but now what do I do with the positive 10

OpenStudy (radar):

Computer went on blink

OpenStudy (anonymous):

haha that's okay

OpenStudy (radar):